# Stochastic Online Conformal Prediction with Semi-Bandit Feedback

### v2

@@ 1 Introduction
Uncertainty quantification is an effective strategy for improving trustworthiness of machine learning models by providing users with a measure of confidence of each prediction to better inform their decisions. Conformal prediction [Vovk et al., 2005, Tibshirani et al., 2019, Park et al., 2019, Bates et al., 2021] has emerged as a promising strategy for uncertainty quantification due to its ability to provide theoretical guarantees for arbitrary blackbox models. They modify a given blackbox model $f:\mathcal{X}\to\mathcal{Y}$ to a conformal predictor $C:\mathcal{X}\to 2^{\mathcal{Y}}$ that predicts sets of labels. In the batch setting, it does so by using a held-out calibration dataset to assess the accuracy of $f$; it constructs smaller prediction sets if $f$ is more accurate and larger ones otherwise. Then, conformal prediction guarantees that the true label is contained in the prediction set with high probability—i.e.,
@@
Traditional conformal prediction require the calibration set to consist of i.i.d. or exchangeable samples from the target distribution; furthermore, the calibration set may need to be large to obtain good performance. In many settings, such labeled data may not be easy to obtain. As a motivating setting, we consider a document retrieval problem, where the input $x$ might be a question and the goal is to retrieve a passage $y$ from a knowledge base such as Wikipedia that can be used to answer $x$; this strategy is known as retrieval-augmented question answering, and can mitigate issues such as hallucinations [Lewis et al., 2020, Shuster et al., 2021, Ji et al., 2023]. A common practice is to use a retrieval model trained on a large dataset in a different domain in a zero-shot manner; for instance, a user might use the dense passage retrieval (DPR) model [Karpukhin et al., 2020] directly on their own dataset without finetuning. Thus, labeled data from the target domain may not be available.
In this paper, we propose an online conformal prediction algorithm that sequentially constructs prediction sets $C_{t}$ as inputs $x_{t}$ are given. We consider semi-bandit feedback, where we only observe the true label $y_{t}^{*}$ if it is included in our prediction set (i.e., $y_{t}^{*}\in C_{t}$). Otherwise, we observe an indicator that $y_{t}^{*}\not\in C_{t}$, but do not observe $y_{t}^{*}$. In the document retrieval example, users might sequentially provide queries $x_{t}$ to the DPR model, and our algorithm responds with a prediction set $C_{t}$ of potentially relevant documents. Then, the user selects the ground truth document $y_{t}^{*}$ if it appears in this set, and indicates that their query was unsuccessful otherwise.
@@
Finally, we empirically evaluate our algorithm on three tasks: image classification, document retrieval, and setting reservation prices in auctions. Our experiments demonstrate that our algorithm generates prediction sets that converge to the optimal ones while maintaining the desired coverage rate. Moreover, our algorithm significantly outperforms three natural baselines; each baseline either achieves worse cumulative expected regret or does not satisfy the desired coverage rate.
@@
Contributions. We formalize and solve the problem of online conformal prediction with semi-bandit feedback. Our algorithm constructs compact prediction sets, ensuring high coverage probability. The algorithm also provides an efficient method for collecting large datasets that are expensive to label. Instead of asking the user to select the ground truth label from the set of all candidate labels, our approach only requires the user to select from a subset. This efficiency gain can be substantial when the label space is large. We assess the algorithm’s performance on image classification, document retrieval and reservation price-setting in auctions, showcasing its effectiveness in a wide variety of real world applications, which highlights the practical utility of our approach.
@@
Related work. Recent work has studied conformal prediction in the online setting [Gibbs and Candes, 2021, Bastani et al., 2022, Gibbs and Candès, 2022, Angelopoulos et al., 2024a, b]. These approaches are motivated by conformal prediction for time series data, where the labels can shift in complex and potentially adversarial ways. As a consequence, they make very different assumptions. In particular, they allow for adversarial rather than i.i.d. assumptions; however, almost all of them assume that the ground truth label is observed at every step, regardless of whether it is contained in the prediction set. For example, the ACI algorithm proposed in Gibbs and Candes [2021] requires observing $y_{t}^{*}$ at every step since it is needed to update the quantile function $\hat{Q}_{t}(\cdot)$, which is needed to compute the prediction set $\hat{C}_{t}(\alpha_{t})$ as well as the loss and gradient update. As the authors point out, they do not need to update the quantile function at every step. However, the steps where they update $\hat{Q}_{t}$ cannot be chosen in a way that depends on $y_{t}^{*}$, since doing so would lead to a biased estimate of the quantile function. In our experiments, we demonstrate that in the semi-bandit feedback setting, updating $\hat{Q}_{t}$ when $y_{t}^{*}$ is in the prediction set can create such a bias that leads ACI to fail to achieve coverage.
@@
Similarly, the SAOCP algorithm proposed by Bhatnagar et al. [2023] requires observing $y^{*}_{t}$ at every step to construct the prediction set $S_{t}=\inf\{s\in R:y^{*}_{t}\in\hat{C}_{t}(X_{t},s)\}$, which is then used to compute the loss and gradient, and the Conformal PID control algorithm proposed by Angelopoulos et al. [2024a] uses $y^{*}_{t}$ to compute the score $s_{t}=s_{t}(x_{t},y^{*}_{t})$, which is needed to compute the gradient. In contrast, we are motivated by the active learning setting, where users are interested in constructing conformal predictors on-the-fly rather than providing a calibration set ahead-of-time. Thus, in our setting, it is reasonable to assume that the data arrives i.i.d.; however, we need to handle semi-bandit feedback since the user may be unable to provide the true label $y_{t}^{*}$ if it is not in the prediction set.
@@
Recently, Angelopoulos et al. [2024b] has pointed out that ACI still works if we take the “quantile function” to be the identity function—i.e., $\hat{Q}_{t}(\alpha)=\alpha$. This strategy achieves the desired coverage rate without requiring observing the ground truth labels $y_{t}^{*}$; instead, they only need to observe whether $y_{t}^{*}$ is contained in the prediction set—i.e., $\mathbbm{1}(y_{t}^{*}\in\hat{C}_{t})$. However, because they forego estimating the quantile function, the resulting algorithm is extremely sensitive to the structure of the scoring function as well as their choice of hyperparameters (in particular, the learning rate). In our experiments, we show that a standard variation the scoring function (in particular, using logits instead of prediction probabilities) causes their performance to significantly degrade.
@@
Angelopoulos et al. [2024b] also point out an additional shortcoming of ACI, which is that it does not guarantee that prediction sets converge to the “optimal” ones in the i.i.d. setting. As noted by Bastani et al. [2022], this issue is reflected in the fact that achieving $\alpha$ coverage (the guarantee satisfied by the ACI algorithm) can be achieved by the “cheating” strategy that outputs $\hat{C}_{t}=\mathcal{Y}$ with probability $1-\alpha$ and $\hat{C}_{t}=\varnothing$ with probability $\alpha$. To remedy this issue, Angelopoulos et al. [2024b] proposes a modification to their algorithm that is guaranteed to converge to the optimal threshold in the setting of i.i.d. observations. However, their result is asymptotic (in addition to suffering from the sensitivity of the scoring function discussed above). Under certain assumptions, we prove that our algorithm satisfies a much stronger regret guarantee; to the best of our knowledge, this is the first regret guarantee on the performance of the prediction sets in terms of convergence of $\tau_{t}$ to $\tau^{*}$.
@@
An additional advantage of our algorithm is that it guarantees that $\tau_{t}\leq\tau^{*}$ with high probability. This guarantee mirrors the distinction between marginal guarantees [Vovk et al., 2005] and probably approximately correct (PAC) (or training conditional) guarantees in the batch setting [Vovk, 2012, Park et al., 2019]. Marginal guarantees have the form $\mathbb{P}_{Z\sim D^{n},(x,y^{*})\sim D}[y^{*}\in C_{Z}(x)]\geq\alpha$, i.e., $\alpha$ coverage over randomness in both the calibration set $Z$ and the new example $(x,y^{*})$. They have the additional advantage that coverage converges to $\alpha$ with the number of calibration examples. In contrast, PAC guarantees disentangle these two sources of randomness:
@@
for a given $\delta\in\mathbb{R}_{>0}$. In other words, coverage holds with high probability over $Z$. Our guarantee that $\tau_{t}\leq\tau^{*}$ is equivalent to providing the coverage guarantee $\mathbb{P}_{(x,y^{*})\sim D}[y^{*}\in C_{\tau_{t}}(x)]\geq\alpha$ for every $t$ with probability at least $1-\delta$ over the whole time horizon. To the best of our knowledge, existing online conformal prediction algorithms all provide a marginal coverage guarantee (it is not even clear how a PAC guarantee would look in the adversarial setting since the calibration examples are not random). For the active learning setting, a PAC guarantee makes sense since it ensures that our algorithm satisfies the desired coverage rate for every user-provided input. Finally, while we do not provide an explicit guarantee that the coverage rate converges to $\alpha$, our regret bound ensures that $\tau_{t}\to\tau^{*}$, which effectively ensures convergence to an $\alpha$ coverage rate.
@@ 2 Problem Formulation
Next, we describe our desired correctness properties. First, given a coverage rate $\alpha\in[0,1)$, our goal is to ensure that we cover the ground truth label with probability at least $\alpha$ on all time steps:
@@
We want (1) to hold with high probability. A trivial solution to the problem described so far is to always take $\tau_{t}=-\infty$. Thus, we additionally want to minimize some measure of the prediction set size. We consider a loss function $\phi:\mathbb{R}\to\mathbb{R}$, where $\phi(\tau_{t})$ encodes the loss incurred on step $t$. Then, our goal is to converge to the best possible prediction sets over time, which we formalize by aiming to achieve sublinear regret. To this end, consider the optimal prediction set $C_{t}^{*}\subseteq\mathcal{Y}$ defined by
@@
$R_{T}=\mathbb{E}\left[\sum_{t\in[T]}\phi(\tau_{t})-\phi(\tau^{*})\right].$
For technical reasons, we need to impose an assumption on our loss. Intuitively, if the loss $\phi$ is discontinuous at $\tau^{*}$, then we may achieve linear regret since $\phi(\tau_{t})-\phi(\tau^{*})\geq c>0$ for some constant $c$. A natural way to formalize the continuity assumption is based on the cumulative distribution function (CDF) of the scoring function. In particular, let $G^{*}$ be the CDF of the random variable $s=f(x,y^{*})$, where $(x,y^{*})\sim D$. Then, we have the following assumption:
@@
That is, if $G^{*}(\tau)$ is very flat in some region, then $\phi(\tau)$ cannot vary very much in that region. Intuitively, if $G^{*}(\tau)$ is flat in some region, then we obtain very few samples in that region, so it becomes very hard to estimate $\tau$ in that region. If $\phi(\tau)$ varies a lot in this region, then the regret can be large if $\tau^{*}$ lies in that region. One caveat is this assumption precludes prediction set size as a loss, since prediction set size is discontinuous. Next, we assume our loss is bounded:
@@
Intuitively, Assumption 2.2 says that the overall reward is bounded, so we cannot accrue huge regret early on in our algorithm when we have very little data. Finally, we make the following assumption:
@@
Note that $G^{*}(\tau)$ is the miscoverage rate of our algorithm for parameter $\tau$. Thus, this assumption says the optimal parameter value $\tau^{*}$ achieves a coverage rate of exactly $\alpha$, which simplifies our analysis.
Then, our goal is to find the optimal prediction sets $C_{t}^{*}$ with coverage rate $\alpha$. Intuitively, $C_{t}^{*}$ is the smallest set that contains the ground truth label with a high probability. At each step, the algorithm observes $x_{t}$ and returns a set $C_{t}$ of candidate labels, and the user either (1) selects the ground truth label $y_{t}^{*}$ from $C_{t}$, or (2) indicates that the ground truth label is not in $C_{t}$. In a document retrieval setting, $x_{t}$ is a query sent by the user, and $y_{t}^{*}$ is the ground truth document, while in an image classification setting, $x_{t}$ is an image and $y_{t}^{*}$ is the ground truth class.
@@ 3 Algorithm
To address this challenge, we instead use a high-probability upper bound on $G^{*}$. In particular, for a error bound $\delta\in\mathbb{R}_{>0}$ to be specified, we construct a $1-\delta$ confidence bound for the empirical CDF $G_{t}$ using the Dvoretzky–Kiefer–Wolfowitz (DKW) inequality [Massart, 1990]. Letting $\overline{G}_{t}$ be the upper confidence bound, we instead aim to choose
@@
Relatedly, our algorithm includes a constraint that $\tau_{t+1}\geq\tau_{t}$ for all $t$. We include this constraint because our estimate $G_{t}$ of the CDF in the interval $[0,\tau_{t})$ may be flawed due to semi-bandit feedback. By avoiding going backwards, we ensure that these flaws do not affect our choice of $\tau_{t+1}$. Again, as long as $\tau_{t}\leq\tau^{*}$, this constraint does not prevent convergence.
$G_{t}(\tau)=\frac{1}{t}\cdot\sum_{j=1}^{t}\mathbbm{1}(\max\{\tau_{t},s_{j}\}\leq\tau).$ (3)
$\overline{G}_{t}(\tau)=G_{t}(\tau)+\epsilon_{t}.$ (4)
We use this $\overline{G}_{t}$ to compute $\tau_{t}$ in Algorithm 1.
Finally, Algorithm 1 satisfies the following theoretical guarantee: (i) it incurs sublinear regret $\tilde{\mathcal{O}}(\sqrt{T})$, and (ii) it satisfies $C_{t}^{*}\subseteq C_{t}$ for all $t$ with probability at least $1-2/T$ (see Appendix A for a proof):
Theorem 3.1.
The expected cumulative regret of Algorithm 1 satisfies
$R_{T}\leq K\left(2\log T+4\sqrt{T\log T}+1\right)+4\phi_{max},$
In addition, with probability at least $1-2/T$, we have $\tau_{t}\leq\tau^{*}$ for all $t$.
@@
     if $s_{t}\geq\tau_{t}$  then
        observe $s_{t}$
     else
        $s_{t}\leftarrow\tau_{t}$
     end if

### v3

@@ 1 Introduction
Uncertainty quantification is an effective strategy for improving trustworthiness of machine learning models by providing users with a measure of confidence of each prediction to better inform their decisions. Conformal prediction (Vovk et al., 2005; Tibshirani et al., 2019; Park et al., 2019; Bates et al., 2021) has emerged as a promising strategy for uncertainty quantification due to its ability to provide theoretical guarantees for arbitrary blackbox models. They modify a given blackbox model $f:\mathcal{X}\to\mathcal{Y}$ to a conformal predictor $C:\mathcal{X}\to 2^{\mathcal{Y}}$ that predicts sets of labels. In the batch setting, it does so by using a held-out calibration dataset to assess the accuracy of $f$; it constructs smaller prediction sets if $f$ is more accurate and larger ones otherwise. Then, conformal prediction guarantees that the true label is contained in the prediction set with high probability—i.e.,
@@
Traditional conformal prediction require the calibration set to consist of i.i.d. or exchangeable samples from the target distribution; furthermore, the calibration set may need to be large to obtain good performance. In many settings, such labeled data may not be easy to obtain. As a motivating setting, we consider a document retrieval problem, where the input $x$ might be a question and the goal is to retrieve a passage $y$ from a knowledge base such as Wikipedia that can be used to answer $x$; this strategy is known as retrieval-augmented question answering, and can mitigate issues such as hallucinations (Lewis et al., 2020; Shuster et al., 2021; Ji et al., 2023). A common practice is to use a retrieval model trained on a large dataset in a different domain in a zero-shot manner; for instance, a user might use the dense passage retrieval (DPR) model (Karpukhin et al., 2020) directly on their own dataset without finetuning. Thus, labeled data from the target domain may not be available.
We propose an online conformal prediction algorithm that sequentially constructs prediction sets $C_{t}$ as inputs $x_{t}$ are given. We consider semi-bandit feedback, where we only observe the true label $y_{t}^{*}$ if it is included in our prediction set (i.e., $y_{t}^{*}\in C_{t}$). Otherwise, we observe an indicator that $y_{t}^{*}\not\in C_{t}$, but do not observe $y_{t}^{*}$. In the document retrieval example, users might sequentially provide queries $x_{t}$ to the DPR model, and our algorithm responds with a prediction set $C_{t}$ of potentially relevant documents. Then, the user selects the ground truth document $y_{t}^{*}$ if it appears in this set, and indicates that their query was unsuccessful otherwise.
@@
We empirically evaluate our algorithm on three tasks: image classification, document retrieval, and setting reservation prices in auctions. Our experiments demonstrate that our algorithm generates prediction sets that converge to the optimal ones while maintaining the desired coverage rate. Moreover, our algorithm significantly outperforms three natural baselines; each baseline either achieves worse cumulative expected regret or does not satisfy the desired coverage rate.
@@
Contributions. We formalize and solve the problem of online conformal prediction with semi-bandit feedback. Our algorithm constructs compact prediction sets, ensuring high coverage probability. The algorithm also provides an efficient method for collecting large datasets that are expensive to label. Instead of asking the user to select the ground truth label from the set of all candidate labels, our approach only requires the user to select from a subset. This efficiency gain can be substantial when the label space is large. We assess our algorithm’s performance on image classification, document retrieval and reservation price-setting in auctions, showcasing its effectiveness in a wide variety of real world applications, which highlights its practical utility.
@@
Related work. Recent work has studied conformal prediction in the online setting (Gibbs & Candes, 2021; Bastani et al., 2022; Gibbs & Candès, 2024; Angelopoulos et al., 2024a, b). They are motivated by conformal prediction for time series data, where the labels can shift in complex and potentially adversarial ways. Thus, they make very different assumptions than ours; in particular, they allow for adversarial rather than i.i.d. assumptions; however, almost all of them assume that the ground truth label is observed at every step, regardless of whether it is contained in the prediction set. For example, the ACI algorithm proposed in Gibbs & Candes (2021) requires observing $y_{t}^{*}$ at every step since it is needed to update the quantile function $\hat{Q}_{t}(\cdot)$, which is needed to compute the prediction set $\hat{C}_{t}(\alpha_{t})$ as well as the loss and gradient update. As the authors point out, they do not need to update the quantile function at every step. However, the steps where they update $\hat{Q}_{t}$ cannot be chosen in a way that depends on $y_{t}^{*}$, since doing so would lead to a biased estimate of the quantile function. In our experiments, we demonstrate that in the semi-bandit feedback setting, updating $\hat{Q}_{t}$ when $y_{t}^{*}$ is in the prediction set can create such a bias that leads ACI to fail to achieve coverage.
@@
Similarly, the SAOCP algorithm proposed by Bhatnagar et al. (2023) requires observing $y^{*}_{t}$ at every step to construct the prediction set $S_{t}=\inf\{s\in R:y^{*}_{t}\in\hat{C}_{t}(X_{t},s)\}$, which is then used to compute the loss and gradient, and the Conformal PID control algorithm proposed by Angelopoulos et al. (2024a) uses $y^{*}_{t}$ to compute the score $s_{t}=s_{t}(x_{t},y^{*}_{t})$, which is needed to compute the gradient. In contrast, we are motivated by the active learning setting, where users are interested in constructing conformal predictors on-the-fly rather than providing a calibration set ahead-of-time. Thus, in our setting, it is reasonable to assume that the data arrives i.i.d.; however, we need to handle semi-bandit feedback since the user may be unable to provide the true label $y_{t}^{*}$ if it is not in the prediction set.
@@
Angelopoulos et al. (2024b) points out that ACI still works if we take the “quantile function” to be the identity function—i.e., $\hat{Q}_{t}(\alpha)=\alpha$. This strategy achieves the desired coverage rate without requiring observing the ground truth labels $y_{t}^{*}$; instead, they only need to observe whether $y_{t}^{*}$ is contained in the prediction set—i.e., $\mathbbm{1}(y_{t}^{*}\in\hat{C}_{t})$. However, because they forego estimating the quantile function, the resulting algorithm is extremely sensitive to the structure of the scoring function as well as their choice of hyperparameters (in particular, the learning rate). In our experiments, we show that a standard variation the scoring function (in particular, using logits instead of prediction probabilities) causes their performance to significantly degrade.
@@
Angelopoulos et al. (2024b) also point out an additional shortcoming of ACI, which is that it does not guarantee that prediction sets converge to the “optimal” ones in the i.i.d. setting. As noted by Bastani et al. (2022), this issue is reflected in the fact that achieving $\alpha$ coverage (the guarantee satisfied by the ACI algorithm) can be achieved by the “cheating” strategy that outputs $\hat{C}_{t}=\mathcal{Y}$ with probability $1-\alpha$ and $\hat{C}_{t}=\varnothing$ with probability $\alpha$. To remedy this issue, Angelopoulos et al. (2024b) proposes a modification to their algorithm that is guaranteed to converge to the optimal threshold in the setting of i.i.d. observations. However, their result is asymptotic (in addition to suffering from the sensitivity of the scoring function discussed above). Under certain assumptions, we prove that our algorithm satisfies a much stronger regret guarantee; to the best of our knowledge, this is the first regret guarantee on the performance of the prediction sets in terms of convergence of $\tau_{t}$ to $\tau^{*}$.
@@
Another advantage of our algorithm is that it guarantees that $\tau_{t}\leq\tau^{*}$ with high probability. This guarantee mirrors the distinction between marginal guarantees (Vovk et al., 2005) and probably approximately correct (PAC) (or training conditional) guarantees (Vovk, 2012; Park et al., 2019). Marginal guarantees have the form $\mathbb{P}_{Z\sim D^{n},(x,y^{*})\sim D}[y^{*}\in C_{Z}(x)]\geq\alpha$, i.e., $\alpha$ coverage over randomness in both the calibration set $Z$ and the new example $(x,y^{*})$. They also have the advantage that coverage converges to $\alpha$ with the number of calibration examples. In contrast, PAC guarantees disentangle these two sources of randomness:
@@
for a given $\delta\in\mathbb{R}_{>0}$. In other words, coverage holds with high probability over $Z$. Our guarantee $\tau_{t}\leq\tau^{*}$ is equivalent to the coverage guarantee $\mathbb{P}_{(x,y^{*})\sim D}[y^{*}\in C_{\tau_{t}}(x)]\geq\alpha$ for every $t$ with probability at least $1-\delta$ over the whole time horizon. To the best of our knowledge, existing online conformal prediction algorithms all provide a marginal coverage guarantee (it is not even clear how a PAC guarantee would look in the adversarial setting since the calibration examples are not random). For the active learning setting, a PAC guarantee makes sense since it ensures our algorithm satisfies the desired coverage rate for every user-provided input. Finally, while we do not provide an explicit guarantee that the coverage converges to $\alpha$, our regret bound ensures that $\tau_{t}\to\tau^{*}$, which ensures coverage converges to $\alpha$.
@@ 2 Problem Formulation
Next, we describe our desired correctness properties. First, given a coverage rate $\alpha\in[0,1)$, our goal is to ensure we cover the true label with probability at least $\alpha$ on all steps:
@@
We want (1) to hold with high probability. A trivial solution to the problem so far is to always take $\tau_{t}=-\infty$. Thus, we additionally want to minimize some measure of prediction set size. We consider a loss function $\phi:\mathbb{R}\to\mathbb{R}$, where $\phi(\tau_{t})$ encodes the loss incurred on step $t$. Then, our goal is to converge to the best possible prediction sets over time, which we formalize by aiming to achieve sublinear regret. Consider the optimal prediction set $C_{t}^{*}\subseteq\mathcal{Y}$ defined by
@@
$R_{T}=\sum_{t\in[T]}\mathbb{E}[\phi(\tau_{t})-\phi(\tau^{*})].$
Intuitively, we need to impose an assumption on our loss because if the loss $\phi$ is discontinuous at $\tau^{*}$, then we may achieve linear regret since $\phi(\tau_{t})-\phi(\tau^{*})\geq c>0$ for some constant $c$. A natural way to formalize the continuity assumption is based on the cumulative distribution function (CDF) of the scoring function. In particular, let $G^{*}$ be the CDF of the random variable $s=f(x,y^{*})$, where $(x,y^{*})\sim D$. Then, we have the following assumption:
@@
That is, if $G^{*}(\tau)$ is very flat in some region, then $\phi(\tau)$ cannot vary much in that region. For example, consider a CDF such that $G^{*}(\tau_{0})=G^{*}(\tau^{*})=1-\alpha$ with $\tau_{0}<\tau^{*}$. Since we never observe any scores between $\tau_{0}$ and $\tau^{*}$, we cannot identify $\tau^{*}$ from $[\tau_{0},\tau^{*}]$ from finitely many samples; even a small error in estimating the empirical CDF (e.g., $G_{t}(\tau^{*})=G^{*}(\tau^{*})+\epsilon$) can result in significant regret.
One caveat is this assumption precludes prediction set size as a loss, since prediction set size is discontinuous. Next, we assume our loss is bounded:
@@
Intuitively, Assumption 2.2 says that the overall reward is bounded, so we cannot accrue huge regret early on in our algorithm when we have very little data. These two assumptions are standard in the bandit literature (Kleinberg et al., 2008). Finally, we make the following assumption:
@@
Note that $G^{*}(\tau)$ is the miscoverage rate of our algorithm for parameter $\tau$. Thus, this assumption says the optimal parameter value $\tau^{*}$ achieves a coverage rate of exactly $\alpha$, which simplifies our analysis. Then, our goal is to find the optimal prediction sets $C_{t}^{*}$ with coverage rate $\alpha$. Intuitively, $C_{t}^{*}$ is the smallest set that contains the ground truth label with a high probability. At each step, the algorithm observes $x_{t}$ and returns a set $C_{t}$ of candidate labels, and the user either (1) selects the ground truth label $y_{t}^{*}$ from $C_{t}$, or (2) indicates that the ground truth label is not in $C_{t}$. In a document retrieval setting, $x_{t}$ is a query sent by the user, and $y_{t}^{*}$ is the ground truth document, while in an image classification setting, $x_{t}$ is an image and $y_{t}^{*}$ is the ground truth class.
@@ 3 Algorithm
To address this challenge, we instead use a high-probability upper bound on $G^{*}$. In particular, for a error bound $\delta\in\mathbb{R}_{>0}$ to be specified, we construct a $1-\delta$ confidence bound for the empirical CDF $G_{t}$ using the Dvoretzky–Kiefer–Wolfowitz (DKW) inequality (Massart, 1990). Letting $\overline{G}_{t}$ be the upper confidence bound, we instead aim to choose
@@
Also, our algorithm includes a constraint $\tau_{t+1}\geq\tau_{t}$ for all $t$. We include this constraint because our estimate $G_{t}$ of the CDF in the interval $[0,\tau_{t})$ may be flawed due to semi-bandit feedback. By avoiding going backwards, we ensure that these flaws do not affect our choice of $\tau_{t+1}$. Again, as long as $\tau_{t}\leq\tau^{*}$, this constraint does not prevent convergence.
@@
     if $s_{t}\geq\tau_{t}$ then observe $s_{t}$ else $s_{t}\leftarrow\tau_{t}$